# SemantiCodec: An Ultra Low Bitrate Semantic Audio Codec for General Sound

### v1

@@ IV-C2 Semantic Information Evaluation
We assess the semantic distinctiveness of the audio codec representation by analyzing classification accuracy on the datasets described in Section IV-A, following the methodology outlined in the HEAR evaluation benchmark [22]. Our evaluation specifically focuses on the quantized output of the codec encoder, which serves as input to the codec decoder. The codec model is frozen to extract the quantized feature without fine-tuning. The feature sequence is averaged along the time axis to obtain the clip-level feature. A shallow downstream multi-layer perceptron (MLP) classifier with two linear layers is trained on clip-level features. The classification performance is reported to indicate the semantic distinctiveness of the codec. Furthermore, several studies on audio language models use powerful AR language models to predict tokens from the first VQ layer while leaving the rest tokens to be predicted by non-AR language models [15]. Therefore, we also report the classification performance using tokens from the first VQ layer to assess the semantic richness.
@@ IV-C3 MUSHRA Test
To assess the subjective audio reconstruction quality of various audio codecs, we adhere to the established MUSHRA test protocol [77]. This method requires participants to evaluate audio samples relative to a concealed reference and predefined anchors, ensuring evaluators are unaware of the original clip to foster impartial assessments. Ratings are assigned on a scale from 00 to $100$. For our MUSHRA evaluation, we randomly select $10$% of the audio from our evaluation set, comprising $25$ music tracks, $30$ speech recordings, and $50$ general sound samples. Participants are tasked with rating the audio quality of nine different samples, which include three variations each of SemanticCodec and Descript codecs at different bitrates, an open-source Encodec at $1.5$kbps, the original audio (ground truth), and an anchor with a low-pass filter applied at $3.5$ kHz. We ensure each set of audio receives evaluations from at least $10$ raters. Participants are instructed with This task evaluates the quality proximity between an audio sample and its reference. Please listen carefully to the reference audio and then rate the quality of each test audio clip compared to the reference. Use the scale where 00 indicates no resemblance to the reference, and $100$ means perfectly the same as the reference. Ratings with ground truth scored below $60$ are excluded to maintain data integrity. Including deliberately degraded anchors allows the MUSHRA test to reveal subtle perceptual distinctions and quality variances across codecs. Our MUSHRA test has received a favourable opinion from the ethics review through completing the University of Surrey self-assessment governance and ethics form. We calculate and report the mean MUSHRA score for each codec as its final subjective evaluation metric.
@@ IV-D Implementation Details
We implemented the k-means clustering on the high-dimensional AudioMAE features extracted from our dataset, as mentioned in Section IV-A. This procedure was carried out individually for three types of audio: music, speech, and general sound. For each type of audio, we run the clustering with different numbers of centroids, including $1024$, $2048$, $4096$, $8192$, and $16384$, to produce $15$ distinct sets of centroids. To manage the computational complexity of clustering high-dimensional features, we made several optimizations to the algorithm, which is available online [^6] https://github.com/haoheliu/kmeans_pytorch.
@@
In building the final semantic codebook for SemantiCodec, we combined codebooks from each of the three audio domains. Given that general audio tends to encompass a broader array of sounds than speech and music, we assigned twice as many centroids. This created four distinct semantic codebooks with varying centroids: $4096$, $8192$, $16384$, and $32768$. For example, the codebook with $16384$ centroids includes $8192$ centroids for general audio and $4096$ centroids each for speech and music. This selection and combination process ensures our SemantiCodec has a comprehensive semantic codebook that accurately captures the diverse range of audio it processes. We repeat the above clustering and merging process for three different settings of stack factors $K\in\{1,2,3\}$. We randomly select and employ one of the four semantic codebooks in each training batch during model optimization. Therefore, our model supports variable vocabulary sizes and bitrates. The acoustic codebook utilizes a fixed-size codebook. We employ $32768$ centroid semantic codebook and $8192$ centroid acoustic codebook by default during model evaluation. Our experiment result indicates that more semantic centroids can moderately improve the reconstruction quality (see Figure 5).
@@
We utilize the AudioMAE model pretrained on AudioSet [^7] https://github.com/facebookresearch/AudioMAE, which features an embedding dimension of $768$. The resulting feature dimension is $K\times 768$ when stacking adjacent frames. Our implementation strategy for both the AudioMAE encoder and the diffusion decoder aligns with the approach described in AudioLDM 2. A notable modification in our setup is the adoption of a more compact T-UNet architecture with a base channel number of $64$ and a parameter number of $75$ million, in contrast to the larger version in AudioLDM 2 which features a base channel number of $128$ with $346$-million parameters. SemantiCodec is configured to operate at three different token rate settings, including $25$, $50$, and $100$ tokens per second. For each setting, the model undergoes training for $500,000$ steps on the designated training set with two A100-Amphere-$80$GB GPUs with a batch size of $48$. With a basic learning rate of $41e-$ 4, we incorporate a linear warm-up phase over the first $5,000$ steps. In the T-UNet architecture, the output from the SemantiCodec encoder, denoted as ${\bm{E}}$, is integrated via cross-attention mechanisms. Given that ${\bm{E}}$ lacks inherent positional information, we enrich it with fixed positional embeddings before input into the cross-attention layers.
@@
SemantiCodec is trained on audio segments exactly $10.24$ seconds in length. To accommodate audio files of varying lengths, we have developed solutions involving either trimming or an overlap-and-add approach. For audio files shorter than $10.24$ seconds, we pad them to the required duration, compute the token encoding, and then remove any tokens corresponding to the padded region. For files longer than $10.24$ seconds, we segment the audio into $10.24$-second chunks without overlap, compute their token embeddings, and concatenate these segments. During inference, decoding uses a window length of $10.24$ seconds with a $6.25\%$ overlap between consecutive windows. The audio content from overlapping segments is blended using a linear decay and gain before combining.
@@ V-A Reconstruction Quality
TABLE I: Objective evaluation of SemantiCodecs and competing baseline codecs at various bitrates on speech, music and general audio. VIS stands for the ViSQOL metric.
General Audio Music Speech Average
@@
Table I shows the audio reconstruction quality of SemantiCodec and baselines described in Section IV-B indicated by objective metrics. We use Token/Sec to denote the number of tokens the audio signal is encoded into per second, which is crucial in audio language modelling as it influences the length of the audio sequence. Given the variability in semantic codebook sizes, models with identical Token/Sec may yield different bit rates (see Figure 5), denoted by kBit/Sec or Kbps. SemantiCodec significantly outperforms the reconstruction quality compared with DAC at similar bitrates. With a low bitrate of $0.71$kbps, the reconstruction quality of SemantiCodec still outperforms the $1.5$kbps Encodec. It is remarkably comparable to HiFi-Codec operating at $2.0$kbps, as the average ViSQOL score indicates. With an ultra-low bitrate of $0.36$kbps, SemantiCodec still achieves a better ViSQOL score than the $1.41$kbps DAC. The $1.43$ kbps SemantiCodec demonstrate performance on par with the $3.0$ kbps Encodec. The superior performance of SemantiCodec in reconstruction at low bitrates indicates its potential in efficient audio transmission and storage and audio-based language modelling since it provides shorter discrete representations of audio without substantially compromising the reconstruction quality.
@@
Using the ground truth unquantized AudioMAE feature as the condition for the latent diffusion model, the setting SemantiCodec w. GT AudioMAE marks the performance upper bound of SemantiCodec. SemantiCodec w. GT AudioMAE demonstrates strong performance on the reconstruction quality with an average ViSQOL score of $4.61$, which indicates that the original AudioMAE feature contains sufficient information to reconstruct the audio signal. However, the metrics MEL and STFT are still considerably lower than other neural codecs, which is expected because SemantiCodec does not directly employ MEL and STFT as the training objective.
@@
By removing the second VQ layer, as shown in the setting of SemantiCodec w.o. Acoustic VQ, the model performance exhibit a considerable degradation, with a WER of $55.6$ and a ViSQOL of $2.61$, highlighting the importance of the second acoustic VQ layer for acoustic reconstruction. Finally, comparing the reconstruction performance across different audio types, it can be concluded that general audio consistently exhibits a slightly inferior reconstruction quality than music and speech. The challenges in reconstructing general audio signals may be attributed to the intrinsic complexity and variability of general audio content. This also validates our choice of assigning more semantic centroid numbers for general audio than speech and music in Section IV-D.
@@
Figure 4: The average MUSHRA test score on our curated evaluation set. Our proposed SemantiCodec outperforms the baseline model even with a significantly lower bitrate.
@@
Figure 4 further showcases the reconstruction performance in terms of subjective MUSHRA scores. We only incorporate codecs with a bitrate of less than or equal to $1.5$kbps into evaluation since we aim at developing codecs with ultra-low bitrates. The result is consistent with the objective evaluation shown in Table I that SemantiCodec significantly outperforms counterparts with a similar or even higher bitrate. With the reference audio being scored to $80.4$, the $1.5$kbps Encodec model achieves a $55.7$ average MUSHRA score, while even our $0.35$ kbps SemantiCodec achieves a significantly higher score of $67.1$. The high MUSHRA score of SemantiCodec can be attributed to the generative nature of the model, which leads to better audio quality at ultra-low bitrate. By comparison, we observe strong artifacts at similar bitrates on DAC and Encodec (shown in Figure 3), which can significantly impact the listener quality.
@@ V-B Semantic in the Codec Tokens
TABLE II: Semantic evaluation of SemantiCodecs and competing baseline codecs using the HEAR benchmark.
VQ setting Model kBit/Sec Token/Sec ESC-50 NSPitch SPC LbCount CRM-D VoImit Average ACC
@@
SemantiCodec $1.43$ $100$ $\mathbf{63.8}$ $73.3$ $\mathbf{43.6}$ $67.0$ $\mathbf{57.9}$ $\mathbf{9.6}$ $\mathbf{52.5}$
$0.71$ $50$ $60.9$ $64.9$ $41.9$ $\mathbf{71.6}$ $53.2$ $9.3$ $50.3$
@@
SemantiCodec $0.71$ $50$ $\mathbf{66.6}$ $\mathbf{73.9}$ $\mathbf{42.7}$ $\mathbf{66.7}$ $\mathbf{57.5}$ $\mathbf{11.1}$ $\mathbf{53.1}$
@@
The semantic richness of different codecs is demonstrated in Table II. First, when all VQ layers are utilized, SemantiCodec significantly outperforms baseline models in semantic information. Notably, even at a low bitrate of $0.35$kbps, the semantic performance of SemantiCodec surpasses that of higher bitrate counterparts like the $6.0$kbps Encodec and DAC. This indicates that AudioMAE features serve as effective means of semantic encoding.
@@
Since many studies utilize tokens from the first VQ layer to train audio language models, we investigate the semantic information in tokens under this situation. As shown in the lower part of Table II, baseline codecs exhibit a substantial drop in semantic performance when switching from full VQ layers to the first layer. For example, the first VQ layer of the $6.0$kbps DAC only achieves an average accuracy of $24.6$. In contrast, SemantiCodec maintains similar semantic accuracy with only the first VQ layer. This supports our assumption that semantic information is primarily encoded by semantic codes obtained through k-means clustering of AudioMAE features, with acoustic details augmented by subsequent acoustic codes. However, there is still a performance gap between unquantized AudioMAE features (ACC $60.6$) and the quantized latent space of SemantiCodec (ACC $53.1$), highlighting the need for further research into mitigating the loss caused by quantization.
@@ V-C Ablation Study
TABLE III: Utilizing a variable semantic codebook size is beneficial for the reconstruction quality, compared with using a fixed semantic codebook with size $32768$.
@@ V-C1 Variable Semantic Codebook Size
As detailed in Section IV-D, during training, we employ a variety of semantic codebooks, enabling SemantiCodec to accommodate different vocabulary sizes. To explore the efficacy of this approach, we conduct experiments comparing a fixed vocabulary size against a variable one, maintaining a semantic codebook size of $32,768$ in both scenarios. We investigate different stack factors $K$ to assess their impact on performance. As presented in Table III, utilizing a variable vocabulary size significantly enhances the average ViSQOL score and reduces the WER. This improvement likely stems from the exposure to diverse codebooks during training, which enhances the ability of the model to interpret and process the quantized AudioMAE features effectively.
TABLE IV: The impact of k-means centroid number on the semantic richness of SemantiCodec.
@@
We also study the effect of semantic codebook size on semantic richness by calculating classification accuracy with the quantized AudioMAE feature. Table IV shows as the centroid number increases, the quantization error induced by k-means modelling is reduced, hence the semantic performance consistently improves. Similar behaviour can be observed in the reconstruction performance, as shown in Figure 5. With more k-means centroids (i.e., semantic vocabulary size) used, reconstructed speech quality steadily improves. However, as shown by Figure 5, higher semantic vocab size can impact the bitrate, posing a trade-off between reconstruction quality and bitrate.

### v2

@@ IV-C2 Semantic Information Evaluation
We assess the semantic distinctiveness of the audio codec representation by analyzing classification accuracy on the datasets described in Section IV-A, following the methodology outlined in the HEAR evaluation benchmark [22]. Our evaluation specifically focuses on the quantized output of the encoder, which serves as input to the decoder. The codec model is frozen to extract the quantized features without fine-tuning. The feature sequences are averaged along the time axis to obtain the clip-level features, which represents the global information of an audio clip. A shallow downstream multi-layer perceptron (MLP) classifier with two linear layers is trained on clip-level features. The classification performance is reported to indicate the semantic distinctiveness of the codec. Furthermore, several studies on audio language models use powerful auto-regressive (AR) language models to predict tokens from the first VQ layer while leaving the rest of the tokens to be predicted by non-AR language models [15]. Therefore, we also report the classification performance using tokens from the first VQ layer to assess the semantic richness.
@@ IV-C3 MUSHRA Test
To assess the subjective audio reconstruction quality of various audio codecs, we adhere to the established MUSHRA test protocol [69]. Participants are required to compare open audio samples against corresponding hidden references, predefined anchors, and the output of different systems. Ratings are assigned on a scale from 00 to $100$. For our MUSHRA evaluation, we randomly select $10$% of the audio from our evaluation set, comprising $25$ music tracks, $30$ speech recordings, and $50$ general sound samples. Participants are tasked with rating the audio quality of nine different samples, which include three variations each of SemanticCodec and Descript codecs at different bitrates, an open-source Encodec at $1.5$ kbps, the original audio (ground truth), and an anchor of the original audio with a low-pass filter applied at $3.5$ kHz. Note that the original MUSHRA test requires two anchor samples with cut-off frequency of $3.5$ kHz and $7$ kHz. However, since our test samples contain frequency information only up to $8$ kHz, the $7$ kHz anchor would not offer sufficient differentiation and was therefore omitted as an anchor in our test. We ensure each set of audio receives evaluations from at least $10$ raters. Participants are instructed with This task evaluates the quality proximity between an audio sample and its reference. Please listen carefully to the reference audio and then rate the quality of each test audio clip compared to the reference. Use the scale where 00 indicates no resemblance to the reference, and $100$ means perfectly the same as the reference. Including deliberately degraded anchors allows the MUSHRA test to reveal subtle perceptual distinctions and quality variances across codecs. We perform post-screening following the standard MUSHRA test protocol [69]. Our MUSHRA test has received a favourable opinion from the ethics review through completing the University of Surrey self-assessment governance and ethics form. We calculate and report the mean MUSHRA score for each codec as its final subjective evaluation metric.
@@ IV-D Implementation Details
We implemented the k-means clustering on the high-dimensional AudioMAE features extracted from our training dataset mentioned in Section IV-A. This procedure was carried out individually for three types of audio: music, speech, and general sound. For each type of audio, we run the clustering with different numbers of centroids, including $1024$, $2048$, $4096$, $8192$, and $16384$, to produce $15$ distinct sets of centroids. To manage the computational complexity of clustering high-dimensional features, we made several optimizations to the algorithm, which is available online [^6] https://github.com/haoheliu/kmeans_pytorch.
@@
In building the final semantic codebook for SemantiCodec, we combined codebooks from each of the three audio domains. Given that general sound tends to encompass a broader array of sounds than speech and music, we assigned twice as many centroids. This created four distinct semantic codebooks with varying number of centroids: $4096$, $8192$, $16384$, and $32768$. For example, the codebook with $16384$ centroids includes $8192$ centroids for general sound and $4096$ centroids each for speech and music. This selection and combination process ensures our SemantiCodec has a comprehensive semantic codebook that accurately captures the diverse range of audio it processes. We repeat the above clustering and merging process for three different settings of stack factors $K\in\{1,2,4\}$. We randomly select and employ one of the four semantic codebooks in each training batch during model optimization. Therefore, our model supports variable vocabulary sizes and bitrates. Table I summarizes the available settings of SemantiCodec with the corresponding codebook sizes, bit rates, and token rates. The acoustic codebook utilizes a fixed-size codebook. We employ $32768$ centroid semantic codebook and $8192$ centroid acoustic codebook by default during model evaluation. Our experiment result indicates that more semantic centroids can moderately improve the reconstruction quality (see Figure 7).
TABLE I: Bit rate allocation table for semantic quantization layer (SEM) and acoustic quantization layer (ACO).
Token/Sec (SEM+ACO) Codebook Size Kbps Overall kbps
SEM ACO SEM ACO
100 $32768$ $8192$ $0.75$ $0.65$ $1.40$
$16384$ $0.70$ $1.35$
$8192$ $0.65$ $1.30$
$4096$ $0.60$ $1.25$
50 $32768$ $8192$ $0.375$ $0.325$ $0.700$
$16384$ $0.350$ $0.675$
$8192$ $0.325$ $0.650$
$4096$ $0.300$ $0.625$
25 $32768$ $8192$ $0.1875$ $0.1625$ $0.3500$
$16384$ $0.1750$ $0.3375$
$8192$ $0.1625$ $0.3250$
$4096$ $0.1500$ $0.3125$
@@
We utilize the AudioMAE model pretrained on AudioSet [^7] https://github.com/facebookresearch/AudioMAE, which features an embedding dimension of $768$. The resulting feature dimension is $K\times 768$ when stacking adjacent frames. Our implementation strategy for both the AudioMAE encoder and the diffusion decoder aligns with the approach described in AudioLDM 2. A notable modification in our setup is the adoption of a more compact T-UNet architecture with a base channel number of $64$ and a parameter number of $75$ million, in contrast to the larger version in AudioLDM 2 which features a base channel number of $128$ with $346$-million parameters. SemantiCodec is configured to operate at three different token rate settings, including $25$, $50$, and $100$ tokens per second. For each setting, the model undergoes training for $500,000$ steps on the designated training set with two A100-Amphere-$80$GB GPUs with a batch size of $48$. With a basic learning rate of $10^{-4}$, we incorporate a linear warm-up phase over the first $5,000$ steps. In the T-UNet architecture, the output from the SemantiCodec encoder, denoted as ${\bm{E}}$, is integrated via cross-attention mechanisms. Given that ${\bm{E}}$ lacks inherent positional information, we enrich it with fixed positional embeddings before input into the cross-attention layers.
@@
SemantiCodec is trained on audio segments exactly $10.24$ seconds in length. To accommodate audio files of varying lengths, we have developed solutions involving either trimming or an overlap-add approach. For audio files shorter than $10.24$ seconds, we pad them to the required duration, compute the token encoding, and then remove any tokens corresponding to the padded region. For files longer than $10.24$ seconds, we segment the audio into $10.24$-second chunks without overlap, compute their token embeddings, and concatenate these segments. During inference, the decoder uses a window length of $10.24$ seconds with a $6.25\%$ overlap between consecutive windows. The audio content from overlapping segments is blended using a linear decay and gain before combining.
@@ V-A Reconstruction Quality
TABLE II: Objective evaluation of SemantiCodecs and competing baseline codecs at various bitrates on speech, music and general sound. VIS stands for the ViSQOL metric.
General Sound Music Speech Average
@@
Table II shows the audio reconstruction quality of SemantiCodec and baselines described in Section IV-B indicated by objective metrics. We use Token/Sec to denote the number of tokens the audio signal is encoded into per second, which is crucial in audio language modelling as it influences the length of the audio sequence. Given the variability in semantic codebook sizes, models with identical token rate may yield different bit rates (see Figure 7), denoted by kbps. SemantiCodec significantly outperforms DAC on the reconstruction quality at similar bitrates. With as low bitrate as $0.70$ kbps, SemantiCodec still outperforms the $1.5$ kbps Encodec on reconstruction quality. It is remarkably comparable to HiFi-Codec operating at $2.0$ kbps, as the average ViSQOL score indicates. With an ultra-low bitrate of $0.36$ kbps, SemantiCodec still achieves a better ViSQOL score than the $1.41$ kbps DAC. The $1.40$ kbps SemantiCodec demonstrates performance on par with the $3.0$ kbps Encodec. The superior performance of SemantiCodec in reconstruction quality at low bitrates indicates its potential in efficient audio transmission, storage and audio-based language modelling since it provides shorter discrete representations of audio without substantially compromising the reconstruction quality.
@@
Using the ground truth unquantized AudioMAE features as the condition for the latent diffusion model, the setting SemantiCodec w. GT AudioMAE marks the performance upper bound of SemantiCodec. SemantiCodec w. GT AudioMAE demonstrates strong performance on the reconstruction quality with an average ViSQOL score of $4.61$, which indicates that the original AudioMAE features contains sufficient information to reconstruct the audio signal. While DAC and Encodec directly use MEL and STFT as training objectives, potentially enhancing their performance on these metrics, SemantiCodec, despite not doing the same, achieves superior results at comparable bitrates, highlighting its overall better performance.
@@
By removing the second VQ layer, as shown in the setting of SemantiCodec w.o. Acoustic VQ, the model performance exhibit a considerable degradation, with a WER of $55.6$ and a ViSQOL of $2.61$, highlighting the importance of the second acoustic VQ layer for acoustic reconstruction. Finally, comparing the reconstruction performance across different audio types, it can be concluded that general sound consistently exhibits a slightly inferior reconstruction quality than music and speech. The challenges in reconstructing general sound signals may be attributed to the intrinsic complexity and variability of general sound content. This also validates our choice of assigning more semantic centroid numbers for general sound than speech and music in Section IV-D.
@@
Figure 3: The average MUSHRA test score on our curated evaluation set. Our proposed SemantiCodec outperforms the baseline models even with a significantly lower bitrate. For fair comparisons, DAC and HiFi-Codec are not included as they use much higher bit rates.
Figure 4: The MUSHRA test score on different domains.
@@
Figure 3 showcases the reconstruction performance in terms of subjective MUSHRA scores. We only incorporate codecs with a bitrate of less than or equal to $1.5$ kbps into evaluation since we aim at developing codecs with ultra-low bitrates. The result is consistent with the objective evaluation shown in Table II that SemantiCodec significantly outperforms counterparts with a similar or even higher bitrate. With the reference audio being scored to $93.9$, the $1.5$ kbps Encodec model achieves a $47.8$ average MUSHRA score, while even our $0.35$ kbps SemantiCodec achieves a significantly higher score of $55.7$. The high MUSHRA score of SemantiCodec can be attributed to the generative nature of the model, which leads to better audio quality at ultra-low bitrate. By comparison, we observe strong artifacts at similar bitrates on DAC and Encodec (shown in Figure 5), which can significantly impact the perceived quality for the listener. Figure 4 provides further analysis of the MUSHRA score on different domains of audio, which shows similar trends as Figure 3. SemantiCodec demonstrates significantly better performance in music, typically involving more complex acoustic information. This suggests that SemantiCodec is highly effective at handling intricate scenarios.
@@ V-B Semantic in the Codec Tokens
TABLE III: Semantic evaluation of SemantiCodecs and competing baseline codecs using the HEAR benchmark.
VQ setting Model Kbps Token/Sec ESC-50 (%) NSPitch (%) SPC (%) LbCount (%) CRM-D (%) VoImit (%) Average (%)
@@
SemantiCodec $1.40$ $100$ $\mathbf{63.8}$ $73.3$ $\mathbf{43.6}$ $67.0$ $\mathbf{57.9}$ $\mathbf{9.6}$ $\mathbf{52.5}$
$0.70$ $50$ $60.9$ $64.9$ $41.9$ $\mathbf{71.6}$ $53.2$ $9.3$ $50.3$
@@
SemantiCodec $0.70$ $50$ $\mathbf{66.6}$ $\mathbf{73.9}$ $\mathbf{42.7}$ $\mathbf{66.7}$ $\mathbf{57.5}$ $\mathbf{11.1}$ $\mathbf{53.1}$
@@
The semantic richness of different codecs is demonstrated in Table III. First, when all VQ layers are utilized, SemantiCodec significantly outperforms baseline models in semantic information. Notably, even at a low bitrate of $0.35$ kbps, the semantic performance of SemantiCodec surpasses that of higher bitrate counterparts like the $6.0$ kbps Encodec and DAC. This indicates that AudioMAE features serve as effective means of semantic encoding.
@@
Since many studies utilize tokens from the first VQ layer to train audio language models, we investigate the semantic information in tokens under this situation. As shown in the lower part of Table III, the baseline codecs exhibit a substantial drop in semantic performance when switching from evaluating all VQ layers to the first layer only. For example, the first VQ layer of the $6.0$ kbps DAC only achieves an average accuracy of $24.6$. In contrast, SemantiCodec maintains similar semantic accuracy with only the first VQ layer. This supports our assumption that semantic information is primarily encoded by semantic codes obtained through k-means clustering of AudioMAE features, with acoustic details augmented by subsequent acoustic codes. However, there is still a performance gap between unquantized AudioMAE features (ACC $60.6$) and the quantized latent space of SemantiCodec (ACC $53.1$), highlighting the need for further research into mitigating the loss caused by quantization.
@@ V-C Ablation Study
TABLE IV: Utilizing a variable semantic codebook size is beneficial for the reconstruction quality, compared with using a fixed semantic codebook of size $32768$.
@@ V-C1 Variable Semantic Codebook Size
As detailed in Section IV-D, during training, we employ a variety of semantic codebooks, enabling SemantiCodec to accommodate different vocabulary sizes. To explore the efficacy of this approach, we conduct experiments comparing a fixed vocabulary size against a variable one, maintaining an acoustic codebook size of $8192$ in both scenarios. We also investigate different stack factors $K$ to assess their impact on performance. As presented in Table IV, utilizing a variable vocabulary size significantly enhances the average ViSQOL score and reduces the WER. This improvement likely stems from the exposure to diverse codebooks during training, which enhances the ability of the model to interpret and process the quantized AudioMAE features effectively.
TABLE V: The impact of k-means centroid number on the semantic richness of SemantiCodec.
@@
We also study the effect of semantic codebook size on semantic richness by calculating classification accuracy with the quantized AudioMAE features. Table V shows as the centroid number increases, the quantization error induced by k-means modelling is reduced, hence the semantic performance consistently improves. Similar behaviour can be observed in the reconstruction performance, as shown in Figure 7. When using a larger number of k-means centroids (i.e., semantic vocabulary size), reconstructed speech quality steadily improves. However, as also shown by Figure 7, higher semantic vocab size will impact the bitrate, posing a trade-off between reconstruction quality and bitrate.